OpenStudy (anonymous):

please help! write an equation of the line satisfying the given conditions. give the answer in slope-intercept form. Parallel to 5x-y=10; y intercept (0,-2)

OpenStudy (debbieg):

What do you know about the slopes of two lines that are parallel?

OpenStudy (anonymous):

they are equal?

OpenStudy (debbieg):

Right! So the slope of the line you are to write the equation for, has the SAME slope as the line given by 5x-y=10. So, do you know how to find the slope of 5x-y=10?

OpenStudy (anonymous):

well, I don't know if it's right but I subtracted 5x and added it to 10. In result, I had -y=10-5x. So I was going to divide by -5 to get x and divide 10 then I got all confused

OpenStudy (***[isuru]***):

yes! the slopes of two functions r equal!! and u can get the first piece of data to ur function from this if the function the question asks has to be in the form of Y = mx + c where m m= gradient and c = intercep Now u already know the gradient or the slope... the slope of the given function is y = 5x -10 sooooo.. the slope is 5 then we can move one step and write our function (the one that problem asks) as Y = 5x + c now all we have to do is find C and that's where the pair of co-ordinates come from the given co - ordinates r ( 0 , -2) which mean the line question asks should pass a point where y = -2 and x = 0 let's substitute those valuse in our equation y = 5x + c -2 = (5x0) + c c = -2 that's it and we r done!! So the function which is parallel to the given function and has the given co-ordinates is y = 5x - 2 hope this will help ya!!!

OpenStudy (debbieg):

LIke @***[ISURU]*** said, you were on the right track - you need to put the given line into the form: y=mx+b Once in that form, you can easily find the slope; it is just the coefficient of the variable x. So you got to this point: -y=10-5x All that is left to get to y=mx + b is to change that coefficient on the y from a -1 to a +1. The easy way to do that is to multiply both sides by -1: (-1)(-y)=(10-5x)(-1) y=-10+5x and then I would probably rearrange that right hand side - not strictly necessary, but it is how we are "used to" seeing the equation for a line: y=5x-10

OpenStudy (anonymous):

ok. thank you. I get where I missed up. How about the next question, It is through the points (4,2) and is perpendicular to x-3y=7

OpenStudy (anonymous):

perpendicular is the opposite isn't it? like it'll be negative and/or positive the equations